# A mathematical theory of topological invariants of quantum spin systems

### v1

@@ 3.1 Observables and derivations
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq\frac{\pi^{8}(n+1)^{4}3^{k}}{1296}\tilde{d}(Z,U)^{-k}.$ (42)
@@
$d(Z,U)$ $\leq d(Z,w)+d(w,U)$ (43)
$\leq d(Z,w)+\operatorname{diam}(X)+d(X,U)$ (44)
$\leq\operatorname{diam}(X\vee Y)+\operatorname{diam}(X)+d(X,U)$ (45)
$\leq 2\operatorname{diam}(X)+\operatorname{diam}(Y)+d(X,U),$ (46)
and so, since (by Lemma 3.3) $\operatorname{diam}(Z)\leq\operatorname{diam}(X\vee Y)\leq\operatorname{diam}(X)+\operatorname{diam}(Y)$, we have
$\tilde{d}(Z,U)$ $\leq 1+3\operatorname{diam}(X)+2\operatorname{diam}(Y)+d(X,U)$ (47)
$\leq 3(1+\operatorname{diam}(X)+d(X,U)+\operatorname{diam}(Y))$ (48)
$\leq 3(1+\operatorname{diam}(X)+d(X,U))(1+\operatorname{diam}(Y))$ (49)
$=3\tilde{d}(X,U)\tilde{d}(Y)$ (50)
where we used the inequality $1+a+b\leq(1+a)(1+b)$ for $a,b\geq 0$. It follows that for any $k>0$ we have
$\tilde{d}(Z,U)^{k}\tilde{d}(X,U)^{-k}\tilde{d}(Y)^{-k}$ $\leq 3^{k}$ (51)
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq 3^{k}\tilde{d}(Z,U)^{-k}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ (52)
@@
$(1+d(z,X)+\operatorname{diam}(X))(1+d(z,Y)+\operatorname{diam}(Y))$ $\leq(1+\operatorname{diam}(X))^{2}(1+\operatorname{diam}(Y))^{2}.$ (53)
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ $\leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}(1+\operatorname{diam}(X))^{-4n-4}(1+\operatorname{diam}(Y))^{-4n-4}$ (54)
$\leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}$ (55)
$\leq\left(\sum_{\begin{subarray}{c}X\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}\right)^{2}$ (56)
$\leq\frac{\pi^{8}(n+1)^{4}}{1296}.$ (57)
@@
$\sum_{X,Y\in{\mathbb{B}}_{n}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$ $=\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$ (58)
$\leq 2\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}$ (59)
$\leq C3^{k}\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\tilde{d}(Z,U)^{-k},$ (60)
with $C=\frac{\pi^{8}(n+1)^{4}}{648}$. This proves that (35) is absolutely convergent and establishes the bound (36).
@@
$\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\left\|\left[{\mathsf{F}}^{X},\left[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}\right]^{Y}\right]^{W}\right\|$ (62)
$\leq$ $4\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\sum_{\begin{subarray}{c}X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Y\subset X^{\prime}\vee Y^{\prime}\end{subarray}}\|{\mathsf{G}}^{X^{\prime}}\|\|{\mathsf{H}}^{Y^{\prime}}\|$ (63)
$\leq$ $C4\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\tilde{d}(Y)^{-4n-4}$ (64)
$\leq$ $C4\|{\mathsf{F}}\|_{{\mathbb{R}}^{n},4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\tilde{d}(X)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ (65)
$<$ $\infty$.∞ . (66)
@@
$[{\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}=\sum_{X,Y,Z\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},[{\mathsf{G}}^{Y},{\mathsf{H}}^{Z}]]^{W}.$ (67)
@@
In particular since $\{0\}$ and ${\mathbb{R}}^{n}$ are transverse, ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ acts on continuously on ${\mathfrak{D}}_{al}(\{0\})$ and this action is easliy seen to extend to a continuous action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathscr{A}}_{a\ell}$ by derivations.
@@ 3.2 Automorphisms
In this section we recall certain automorphisms obtained by exponentiating elements of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ following [11]. One can develop the theory of such automorphisms that are almost-localized on regions in ${\mathbb{R}}^{n}$ in a similar spirit to the above, but since we do not have much occasion to use them in this work, we opt instead for a more minimal development. Let ${\mathsf{F}}:{\mathbb{R}}\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ be a smooth map. It is shown in [11] that for any $U\subset{\mathbb{R}}^{n}$ and any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}(U)$ the differential equation
$\frac{d}{dt}{\mathcal{A}}(t)=[{\mathsf{F}}(t),{\mathcal{A}}(t)]$ (68)
with the initial condition ${\mathcal{A}}(0)={\mathcal{A}}$ has a unique solution ${\mathcal{A}}(t)\in{\mathscr{A}}_{a\ell}(U)$ for all $t\in{\mathbb{R}}$. Denote by $\alpha^{\mathsf{F}}_{t}$ the map taking ${\mathcal{A}}$ to ${\mathcal{A}}(t)$. It is a continuous automorphism of the Lie algebra ${\mathscr{A}}_{a\ell}(U)$ that preserves the *-operation.
@@
We call any $\alpha^{\mathsf{F}}_{t}$ obtained as above a locally-generated automorphism or LGA for short.
It is further shown in [11] that every LGA extends to a continuous *-automorphism of the quasilocal algebra ${\mathscr{A}}$, and that the set of LGAs forms a group under composition.
@@ 3.3 States
By a state $\psi$ we will mean a state on the quasilocal algebra ${\mathscr{A}}$. If ${\mathsf{F}}$ is an inner derivation (see Def. 3.3), we define its $\psi$-average as the evaluation of $\psi$ on the corresponding element of ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}_{a\ell}$. The group of LGAs acts on states by precomposition, which we denote $\psi^{\alpha}:=\psi\circ\alpha$. We say an LGA $\alpha$ preserves a state $\psi$ if $\psi^{\alpha}=\psi$. We say an element ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ preserves $\psi$ if $\psi([{\mathsf{F}},{\mathcal{A}}])=0$ for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}(\{0\})$, which is equivalent to the one-parameter group $\alpha_{t}^{\mathsf{F}}$ preserving $\psi$.
@@
It is easy to check that ${\mathfrak{D}}_{al}^{\psi}(U)$ is a closed subset of ${\mathscr{A}}_{a\ell}(U)$, and that if ${\mathcal{A}}$ and ${\mathcal{B}}$ preserve $\psi$ then $[{\mathcal{A}},{\mathcal{B}}]$ preserves $\psi$, and as a result the analog of Propositions 3.4 and 3.6 and Corollary 3.1 hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$. Proposition 3.5 on the other hand, does not easily extend to the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$. To do so, we will restrict to gapped states, where the quasiadiabatic generator can be used to prove the analog of Proposition 3.5.
@@
A state $\psi$ is gapped if there exists ${\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ and $\Delta>0$ such that for any ${\mathcal{A}}\in{\mathscr{A}}$ one has
@@
The meaning of this condition becomes more transparent if one recalls that any ${\mathsf{H}}\in{\mathfrak{D}}_{al}$ is a generator of a 1-parameter group of *-automorphisms of ${\mathscr{A}}_{a\ell}$ [11]. The condition (69) implies that $\psi$ is invariant under this 1-parameter group of automorphisms [18], and that the corresponding 1-parameter group of unitaries in the GNS representation of ${\mathscr{A}}$ has a generator whose spectrum in the orthogonal complement to the GNS vacuum vector is contained in $[\Delta,+\infty)$. The condition (69) also implies that $\psi$ is pure [19].
@@
If $\psi$ is a gapped state of ${\mathscr{A}}$ and $\alpha$ is an LGA, then $\psi\circ\alpha$ is also a gapped state. In [11] it was proposed to define a gapped phase as an orbit of gapped state under the action of the group of LGAs.
@@
$\mathcal{J}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}^{\psi}({\mathbb{R}}^{n})$ (70)
$\mathcal{K}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ (71)
@@
    $\|\mathcal{J}({\mathsf{F}})\|_{U,k}$ $\leq C_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$ (72)
    $\|\mathcal{K}({\mathsf{F}})\|_{U,k}$ $\leq C^{\prime}_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$ (73)
@@
    ${\mathsf{F}}=\mathcal{J}({\mathsf{F}})-\mathcal{K}([{\mathsf{H}},{\mathsf{F}}]).$ (74)
@@
    If $\psi$ is gapped then the sequence (30) admits a right splitting, and in particular it is right-exact.
@@
    If $U$ and $V$ are transverse, then (30) is left-exact.
@@
The proof of Proposition 3.5 goes through unmodified except for the definition of $\gamma$, which needs to be changed to ensure that the image of $\gamma$ consists of operators that preserve $\psi$. Suppose $U,V\subseteq{\mathbb{R}}^{n}$ and ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U\cup V)$. Define
$U^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\}$ (82)
$V^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,V)\leq d(x,U)\}$ (83)
@@
$\tilde{\gamma}_{i}({\mathsf{F}})$ $:=\mathcal{J}(\gamma_{i}^{U,V}({\mathsf{F}}))-\mathcal{K}([\mathcal{J}(\gamma_{i}^{U^{\prime},V^{\prime}}({\mathsf{H}})),{\mathsf{F}}])$ (84)

### v2

@@ 3.1 Observables and derivations
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq\frac{\pi^{8}16^{n}(n+1)^{4}3^{k}}{1296}\tilde{d}(Z,U)^{-k}.$
@@
$d(Z,U)$ $\leq d(Z,w)+d(w,U)$
$\leq d(Z,w)+\operatorname{diam}(X)+d(X,U)$
$\leq\operatorname{diam}(X\vee Y)+\operatorname{diam}(X)+d(X,U)$
$\leq 2\operatorname{diam}(X)+\operatorname{diam}(Y)+d(X,U),$
and so, since by Lemma 3.3 $\operatorname{diam}(Z)\leq\operatorname{diam}(X\vee Y)\leq\operatorname{diam}(X)+\operatorname{diam}(Y)$, we have
$\tilde{d}(Z,U)$ $\leq 1+3\operatorname{diam}(X)+2\operatorname{diam}(Y)+d(X,U)$
$\leq 3(1+\operatorname{diam}(X)+d(X,U)+\operatorname{diam}(Y))$
$\leq 3(1+\operatorname{diam}(X)+d(X,U))(1+\operatorname{diam}(Y))$
$=3\tilde{d}(X,U)\tilde{d}(Y).$ (14)
It follows that for any $k>0$ we have
$\tilde{d}(Z,U)^{k}\tilde{d}(X,U)^{-k}\tilde{d}(Y)^{-k}$ $\leq 3^{k}$
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq$
$\hskip 28.45274pt\leq 3^{k}\tilde{d}(Z,U)^{-k}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ (15)
@@
$(1+d(z,X)+\operatorname{diam}(X))(1+d(z,Y)+\operatorname{diam}(Y))\leq$
$\hskip 42.67912pt\leq(1+\operatorname{diam}(X))^{2}(1+\operatorname{diam}(Y))^{2}.$ (16)
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}\leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}(1+\operatorname{diam}(X))^{-4n-4}(1+\operatorname{diam}(Y))^{-4n-4}\\ \leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}\\ \leq\left(\sum_{\begin{subarray}{c}X\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}\right)^{2}\leq\frac{\pi^{8}16^{n}(n+1)^{4}}{1296}.$ (17)
@@
$\sum_{X,Y\in{\mathbb{B}}_{n}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$ $=\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$
$\leq 2\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}$
$\leq C3^{k}\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\tilde{d}(Z,U)^{-k},$
with $C=\frac{\pi^{8}16^{n}(n+1)^{4}}{648}$. This proves that (11) is absolutely convergent and establishes the bound (12).
@@
$\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\left\|\left[{\mathsf{F}}^{X},\left[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}\right]^{Y}\right]^{W}\right\|$
$\leq$ $4\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\sum_{\begin{subarray}{c}X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Y\subset X^{\prime}\vee Y^{\prime}\end{subarray}}\|{\mathsf{G}}^{X^{\prime}}\|\|{\mathsf{H}}^{Y^{\prime}}\|$
$\leq$ $C4\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\tilde{d}(Y)^{-4n-4}$
$\leq$ $C4\|{\mathsf{F}}\|_{{\mathbb{R}}^{n},4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\tilde{d}(X)^{-4n-4}\tilde{d}(Y)^{-4n-4}$
$<$ $\infty$.∞ .
@@
$[{\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}=\sum_{X,Y,Z\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},[{\mathsf{G}}^{Y},{\mathsf{H}}^{Z}]]^{W}.$
@@
In particular since $\{0\}$ and ${\mathbb{R}}^{n}$ are transverse, ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ acts continuously on ${\mathfrak{D}}_{al}(\{0\})$ and this action is easily seen to extend to a continuous action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on the space ${\mathscr{A}}_{a\ell}$. We denote the action of ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ by ${\mathcal{A}}\mapsto{\mathsf{F}}({\mathcal{A}})$. By Prop. 3.6, it is given by
${\mathsf{F}}({\mathcal{A}})=\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}].$ (19)
It is not hard to check that for any $X\in{\mathbb{B}}_{n}$ and any ${\mathcal{A}}\in{\mathscr{A}}(X)$ we have ${\overline{\rm tr}}_{X^{c}}\left({\mathsf{F}}({\mathcal{A}})\right)=[{\mathsf{F}}_{X},{\mathcal{A}}]$ and so the action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathscr{A}}_{a\ell}$ is faithful. Thus, elements of ${\mathfrak{D}}_{al}(U)$ for any $U\subset{\mathbb{R}}^{n}$ may be identified with a subset of the Fréchet-continuous derivations of ${\mathscr{A}}_{a\ell}$. By Proposition 3.2, the Fréchet-Lie algebra ${\mathfrak{D}}_{al}(\{0\})$ is identified with the Fréchet-Lie algebra ${{\mathfrak{d}}_{al}}$ of traceless anti-hermitian elements of ${\mathscr{A}}_{a\ell}$ acting by inner derivations.
@@ 3.2 Automorphisms
In this section we recall certain automorphisms obtained by exponentiating elements of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ following [11]. One can develop the theory of such automorphisms that are almost-localized on regions in ${\mathbb{R}}^{n}$ in a similar spirit to the above, but since we do not have much occasion to use them in this work, we opt instead for a more minimal development. Let ${\mathsf{F}}:{\mathbb{R}}\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ be a smooth map. It is shown in [11] that for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ the differential equation
$\frac{d}{dt}{\mathcal{A}}(t)={\mathsf{F}}(t)({\mathcal{A}}(t))$
with the initial condition ${\mathcal{A}}(0)={\mathcal{A}}$ has a unique solution ${\mathcal{A}}(t)\in{\mathscr{A}}_{a\ell}$ for all $t\in{\mathbb{R}}$. Denote by $\alpha^{\mathsf{F}}_{t}$ the map taking ${\mathcal{A}}$ to ${\mathcal{A}}(t)$. It is a continuous automorphism of the Lie algebra ${\mathscr{A}}_{a\ell}$ that preserves the *-operation.
@@
We call any automorphism of the form $\alpha^{\mathsf{F}}_{t}$ for some smooth map ${\mathsf{F}}:{\mathbb{R}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ a locally-generated automorphism, or LGA for short.
It is shown in [11] that every LGA extends to a continuous *-automorphism of the quasilocal algebra ${\mathscr{A}}$, and that the set of LGAs forms a group under composition.
@@ 3.3 States
By a state $\psi$ we will mean a state on the quasilocal algebra ${\mathscr{A}}$. If ${\mathsf{F}}$ is an inner derivation (see Def. 3.4), we define its $\psi$-average as the evaluation of $\psi$ on the corresponding element of ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}_{a\ell}$. The group of LGAs acts on states by pre-composition, which we denote $\psi^{\alpha}:=\psi\circ\alpha$. We say an LGA $\alpha$ preserves a state $\psi$ if $\psi^{\alpha}=\psi$. We say an element ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ preserves $\psi$ if $\psi({\mathsf{F}}({\mathcal{A}}))=0$ for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$, which is equivalent to the one-parameter group of automorphisms $t\mapsto\alpha_{t}^{\mathsf{F}}$ corresponding to a constant map $t\mapsto{\mathsf{F}}$ preserving $\psi$.
@@
It is easy to check that ${\mathfrak{D}}_{al}^{\psi}(U)$ is a closed subset of ${\mathfrak{D}}_{al}(U)$, and that if ${\mathsf{F}}$ and ${\mathsf{G}}$ preserve $\psi$, then $[{\mathsf{F}},{\mathsf{G}}]$ preserves $\psi$. Thus the analog of Propositions 3.4 and 3.6 and Corollary 3.1 hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$. Proposition 3.5 on the other hand, does not hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$ for a general state $\psi$. To circumvent this, we will restrict to gapped states, where quasiadiabatic evolution [18, 19, 20] can be used to prove the analog of Proposition 3.5.
@@
A state $\psi$ is gapped if there exists ${\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ and $\Delta>0$ such that for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ one has
@@
The meaning of this condition becomes more transparent if one recalls that any ${\mathsf{H}}\in{\mathfrak{D}}_{al}$ is a generator of a one-parameter group of *-automorphisms of ${\mathscr{A}}$ [11]. The condition (20) implies that $\psi$ is invariant under this one-parameter group of automorphisms [21], and that the corresponding one-parameter group of unitaries in the GNS representation of ${\mathscr{A}}$ has a generator whose spectrum in the orthogonal complement to the GNS vacuum vector is contained in $[\Delta,+\infty)$. The condition (20) also implies that $\psi$ is pure [22].
@@
If $\psi$ is a gapped state of ${\mathscr{A}}$ and $\alpha$ is an LGA, then $\psi^{\alpha}$ is also a gapped state. In [11] it was proposed to define a gapped phase as an orbit of gapped state under the action of the group of LGAs.
@@
$\mathcal{J}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}^{\psi}({\mathbb{R}}^{n})$
$\mathcal{K}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$
@@
    $\|\mathcal{J}({\mathsf{F}})\|_{U,k}$ $\leq C_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$
    $\|\mathcal{K}({\mathsf{F}})\|_{U,k}$ $\leq C^{\prime}_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$
@@
    ${\mathsf{F}}=\mathcal{J}({\mathsf{F}})-\mathcal{K}([{\mathsf{H}},{\mathsf{F}}]).$
@@
    If $\psi$ is gapped then the sequence (10) admits a right splitting, and in particular it is exact on the right
@@
    If $U$ and $V$ are transverse, then (10) is exact on the left.
@@
The proof of Proposition 3.5 goes through unmodified except for the definition of $\gamma$, which needs to be changed to ensure that the image of $\gamma$ consists of derivations that preserve $\psi$. Suppose $U,V\subseteq{\mathbb{R}}^{n}$ and ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U\cup V)$. Define
$U^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\},$
$V^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,V)\leq d(x,U)\}.$
@@
$\tilde{\gamma}_{i}({\mathsf{F}})$ $:=\mathcal{J}(\gamma_{i}^{U,V}({\mathsf{F}}))-\mathcal{K}([\mathcal{J}(\gamma_{i}^{U^{\prime},V^{\prime}}({\mathsf{H}})),{\mathsf{F}}])$